OpenStudy (goformit100):

For the equation only :- Sachin scored twice as many runs as Rahul. Together, their runs fell two short of a double century. How many runs did each one score ?

OpenStudy (anonymous):

double century? what poet wrote this question?

OpenStudy (goformit100):

it is a question from Indian writer.

OpenStudy (campbell_st):

rahul = x sachin = 2x so x + 2x = 198 solve for x

OpenStudy (anonymous):

i would say \[2S+S=198\]

OpenStudy (anonymous):

Rahul = x Sachin = 2x x+2x=198

OpenStudy (anonymous):

or if you are a fan of \(x\) then either answer above

OpenStudy (anonymous):

lool

OpenStudy (goformit100):

How youll wrote 198 ?

OpenStudy (anonymous):

lol is right, i am wrong

OpenStudy (anonymous):

It says fell two short of a double century.